# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@ 2.1. Equivariant sectional category
Suppose $T$ is the trivial subgroup of $G$. By [20, Theorem 4], we know that for a $G$-fibration $p$, the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is fibration for all closed subgroups $H$ of $G$. Hence, by taking $K=T$ in Proposition 2.7, it follows that $\mathrm{secat}(p^{H})=\mathrm{secat}_{T}(p^{H})\leq\mathrm{secat}_{G}(p)$.
By [20, Theorem 3], we know that $p$ is a $K$-fibration for all closed subgroups $K$ of $G$. Hence, by taking $H=T$ in Proposition 2.7, we get the subgroup inequality $\mathrm{secat}_{K}(p)=\mathrm{secat}_{K}(p^{T})\leq\mathrm{secat}_{G}(p)$. Note that $T$ is a closed subgroup of compact Hausdorff topological group $K$. Hence, applying the subgroup inequality for the $K$-fibration $p$, we get $\mathrm{secat}(p)=\mathrm{secat}_{T}(p)\leq\mathrm{secat}_{K}(p)$. ∎
The following proposition states some basic properties of the equivariant sectional category. Proofs are left to the reader. For analogous results concerning the non-equivariant sectional category, refer to [22, Lemma 2.1].
@@
by Proposition 2.9. ∎
Generalizing Schwarz’s dimension-connectivity upper bound on the sectional category, Grant established the corresponding equivariant analogue for the equivariant sectional category in [21, Theorem 3.5]. We extend this approach to derive an equivariant homotopy dimension-connectivity upper bound for equivariant sectional category. To achieve this, we first introduce the notion of $G$-homotopy dimension for $G$-$\mathrm{CW}$-complexes.
@@
Suppose $p\colon E\to B$ is a Serre $G$-fibration with fibre $F$, whose base $B$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$. If there exists $s\geq 0$ such that the fibre of $p^{H}\colon E^{H}\to B^{H}$ is $(s-1)$-connected for all subgroups $H$ of $G$, then
@@
It is enough to show that for any $G$-$\mathrm{CW}$-complex $B^{\prime}$ which is $G$-homotopy equivalent to $B$, we have
$\mathrm{secat}_{G}(p)<\frac{\mathrm{dim(B^{\prime})+1}}{s+1}+1.$
Suppose $f\colon B^{\prime}\to B$ is a $G$-homotopy between $G$-$\mathrm{CW}$-complexes $B^{\prime}$ and $B$, and $p^{\prime}\colon E^{\prime}\to B^{\prime}$ is the pullback of $p$ along $f$. Then, by Corollary 2.10, we have $\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(p)$. Since the fibre of $p^{\prime}$ is also $F$, we get
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is a compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts on $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$ diagonally. Furthermore, if $B_{1}$ and $B_{2}$ are Hausdorff, and $B_{1}\times B_{2}$ is completely normal, then
@@
Suppose $G$ is compact Hausdorff. Then identifying $G$ with the diagonal subgroup of $G\times G$, we see that it is a closed subgroup of $G\times G$. Hence, by [20, Theorem 3], it follows that $p_{1}\times p_{2}\colon E_{2}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$.
If $B_{1}$ and $B_{2}$ are Hausdorff, and $B_{1}\times B_{2}$ is completely normal, then [10, Lemma 3.12] implies $B_{1}\times B_{2}$ is $(G\times G)$-completely normal. Hence, the desired inequality
@@
follows from Corollary 2.8 (2) and [1, Proposition 3.7]. ∎
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $B$ is Hausdorff and $B\times B$ is completely normal, then
@@
[…]
is a pullback in the category of $G$-spaces, where $\Delta\colon B\to B\times B$ is the diagonal map. In Proposition 2.14, we showed that $p_{1}\times p_{2}$ is a $G$-fibration if $G$ is compact Hausdorff. Hence, $p$ is a $G$-fibration. Thus, the desired inequality
@@
follows from [10, Proposition 4.3] and Proposition 2.14. ∎
@@ 2.2. Equivariant LS-category
The notion of Lusternik-Schnirelmann (LS) category was introduced by Lusternik and Schnirelmann in [26]. In this section, we present the corresponding equivariant analogue.
@@
Definition 2.17.
[14] The equivariant LS-category of a $G$-space $X$, denoted by $\mathrm{cat}_{G}(X)$, is the least positive integer $k$ such that there exists a $G$-categorical open cover $\{U_{1},\dots,U_{k}\}$ of $X$.
@@
Lemma 2.19 ([10, Corollary 4.7]).
If $X$ is a $G$-space such that $X$ is $G$-connected and $x_{0}\in X^{G}$, then $\mathrm{cat}_{G}(X)=\mathrm{secat}_{G}(e_{X}).$
@@
Suppose $X$ is a $G$-connected space with $X^{G}\neq\emptyset$. If $H$ and $K$ are closed subgroups of $X$ such that $X^{H}$ is $K$-invariant, then
@@
In particular, if $G$ is Hausdorff, then
@@
Proof.
We note that $X^{H}$ is $K$-connected since $X$ is $G$-connected, and $H$ and $K$ are closed subgroups of $G$. If $x_{0}\in X^{G}$, then $x_{0}\in(X^{H})^{K}=X^{H\cap K}$. Hence, by Lemma 2.19, it is enough to show that $\mathrm{secat}_{K}(e_{X^{H}})\leq\mathrm{secat}_{G}(e_{X})$.
Now, as a consequence of Theorem 2.12, we obtain an equivariant homotopy dimension-connectivity upper bound for equivariant LS category.
@@
Suppose $X$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$ such that $X^{G}\neq\emptyset$. If there exists $s\geq 0$ such that $X^{H}$ is $s$-connected for all subgroups $H$ of $G$, then
@@
If $x_{0}\in X^{G}$, then $e_{X}\colon P_{x_{0}}X\to X$ is a $G$-fibration with fibre $\Omega X$ which also admits a $G$-action. Note that $(\Omega X)^{H}=\Omega(X^{H})$. Since $X^{H}$ is $s$-connected, the loop space $\Omega(X^{H})$ is $(s-1)$-connected. Hence, by Theorem 2.12, we get
@@
As $X^{H}$ is $s$-connected, it follows that $X^{H}$ is path-connected. Hence, $X$ is $G$-connected, and the theorem follows by Lemma 2.19. ∎
@@ 2.3. Equivariant and invariant topological complexity
We recall the concept of equivariant topological complexity introduced by Colman and Grant in [10]. Let $X$ be a $G$-space. Observe that the free path space $PX$ admits a $G$-action via $(g\cdot\alpha)(t):=g\cdot\alpha(t)$. Similarly, the product space $X^{k}$ is a $G$-space with the diagonal action. The fibration
@@
In [1, Proposition 3.40], Sarkar and the authors of this paper provided a dimension-connectivity upper bound on the sequential equivariant topological complexity. We improve their result by establishing an equivariant homotopy dimension-connectivity upper bound. We omit the proof, as it is similar to the original and follows from the homotopy dimension-connectivity upper bound on the equivariant sectional category in Theorem 2.12.
@@
Let $X$ be a $G$-$\mathrm{CW}$-complex of dimension atleast 1 such that $X^{H}$ is $m$-connected for all subgroups $H\leq G$. Then
$\mathrm{TC}_{k,G}(X)<\frac{k\ \mathrm{hdim}_{G}(X)+1}{m+1}+1.$
It is important to note that the equivariant topological complexity of $G$-spaces does not necessarily relate to the topological complexity of their orbit spaces. However, Lubawski and Marzantowicz provided an alternative definition of equivariant topological complexity, designed to facilitate such a comparison. We now present their definition and recall the corresponding result.
Suppose $X$ is a $G$-space. Let $\pi_{X}:X\to X/G$ denote the orbit map.
@@
That is the following diagram
${PX\times_{X/G}PX}$${PX}$${PX}$${X/G}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{X}\circ e_{0}}$$\scriptstyle{\pi_{X}\circ e_{1}}$
@@
$\mathfrak{p}\colon PX\times_{X/G}PX\to X\times X,\quad(\gamma,\delta)\mapsto(\gamma(0),\delta(1)).$
It was shown in [25, Proposition 3.7] that the map $\mathfrak{p}$ is a $(G\times G)$-fibration.
@@
Moreover, if $X$ has one orbit type, then
@@ 2.4. Clapp-Puppe invariant of Lusternik-Schnirelmann type
Colman and Grant in [10, Lemma 5.14] showed that for a $G$-invariant open subset $U$ of $X\times X$ the following are equivalent:
@@
Later, Lubawski and Marzantowicz in [25, Lemma 3.8] showed a similar result for invariant topological complexity. More precisely, for a $(G\times G)$-invariant open subset of $U$ of $X\times X$ the following are equivalent:
@@
In Section 3 and Section 4, we give analogous results for equivariant parametrized topological complexity and invariant parametrized topological complexity, respectively. We use these results to prove Theorem 4.29.

### v4

@@ 2.2. Equivariant sectional category
If $s\colon U\to E$ is a $G$-equivariant section of $p$, then $\left.s\right|_{V}\colon V\to E^{H}$ is a $K$-equivariant section of $p^{H}$, where $V=U\cap B^{H}$. Hence, the inequality (4) follows.
If $G$ is compact Hausdorff, then it follows from [gevorgyan2023equivariant, Theorem 4 and Theorem 3] that $p^{H}\colon E^{H}\to B^{H}$ and $p\colon E\to B$ are a fibration and a $K$-fibration, respectively. Hence, the following inequalities follow by taking $K$ and $H$ to be the trivial subgroup, respectively. ∎
The following proposition states some basic properties of the equivariant sectional category. Proofs are left to the reader. For analogous results concerning the non-equivariant sectional category, we refer to [Grant-ParaTC-group, Lemma 2.1].
@@
by Proposition˜2.13. ∎
Generalizing Schwarz’s dimension-connectivity upper bound on the sectional category, Grant established the corresponding equivariant analogue for the equivariant sectional category in [grant2019symmetrized, Theorem 3.5]. We extend this approach to derive an equivariant homotopy dimension-connectivity upper bound for equivariant sectional category. To achieve this, we first introduce the notion of $G$-homotopy dimension for $G$-$\mathrm{CW}$-complexes.
@@
Suppose $G$ is a compact Lie group. Suppose $p\colon E\to B$ is a $G$-fibration with fibre $F$, whose base $B$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$. If there exists $s\geq 0$ such that the fibre of $p^{H}\colon E^{H}\to B^{H}$ is (s1)(s-1)-connected for all subgroups $H$ of $G$, then
@@
Suppose $f\colon B^{\prime}\to B$ is a $G$-homotopy between $G$-$\mathrm{CW}$-complexes $B^{\prime}$ and $B$, and $p^{\prime}\colon E^{\prime}\to B^{\prime}$ is the pullback of $p$ along $f$. Then, by Corollary˜2.14, we have $\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(p)$. Since the fibre of $p^{\prime}$ is also $F$, we get
@@
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts on $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$ diagonally. Furthermore, if $B_{1}\times B_{2}$ is completely normal, then
@@
Identifying $G$ with the diagonal subgroup of $G\times G$, we see that it is a closed subgroup of $G\times G$. Hence, by [gevorgyan2023equivariant, Theorem 3], it follows that $p_{1}\times p_{2}$ is a $G$-fibration. Furthermore, by [colmangranteqtc, Lemma 3.12], it follows that $B_{1}\times B_{2}$ is $(G\times G)$-completely normal. Hence, the desired inequality
@@
follows from Proposition˜2.12 (2) and [A-D-S, Proposition 3.7]. ∎
@@
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $B\times B$ is completely normal, then
@@
$E_{1}\times_{B}E_{2}$$E_{1}\times E_{2}$$B{B\times B}$$\scriptstyle{p}$$\scriptstyle{p_{1}\times p_{2}}$$\scriptstyle{\Delta}$ (5)
is a pullback in the category of $G$-spaces, where $\Delta\colon B\to B\times B$ is the diagonal map. In Proposition˜2.18, we showed that $p_{1}\times p_{2}$ is a $G$-fibration if $G$ is compact Hausdorff. Hence, $p$ is a $G$-fibration. Thus, the desired inequality
@@
follows from [colmangranteqtc, Proposition 4.3] and Proposition˜2.18. ∎
@@ 2.3. Equivariant LS-category
The notion of Lusternik-Schnirelmann (LS) category was introduced by Lusternik and Schnirelmann in [LScat]. In this section, we recall the corresponding equivariant analogue.
@@
Definition 2.21 ([Fadelleqcat]).
The equivariant LS-category of a $G$-space $X$, denoted by $\mathrm{cat}_{G}(X)$, is the least positive integer $k$ such that there exists a $G$-categorical open cover $\{U_{1},\dots,U_{k}\}$ of $X$.
@@
Lemma 2.23 ([colmangranteqtc, Corollary 4.7]).
Suppose $G$ is a compact Hausdorff topological group. If $X$ is a $G$-space such that $X$ is $G$-connected and $x_{0}\in X^{G}$, then $\mathrm{cat}_{G}(X)=\mathrm{secat}_{G}(e_{X}).$
@@
Let $G$ be a compact Hausdorff topological group, and let $X$ be a $G$-connected space with $X^{G}\neq\emptyset$. If $H$ and $K$ are closed subgroups of $G$ such that $X^{H}$ is $K$-invariant and $HK^{\prime}=K^{\prime}H$ for all closed subgroups $K^{\prime}$ of $K$, then
@@
Proof.
Suppose $K^{\prime}$ is a closed subgroup of $K$. Then $HK^{\prime}=K^{\prime}H$ implies $HK^{\prime}$ is a subgroup of $G$ and $\langle H,K^{\prime}\rangle=HK^{\prime}$, where $\langle H,K^{\prime}\rangle$ is the subgroup generated by $H$ and $K^{\prime}$. We note that $HK^{\prime}$ is closed, as it is the image of the compact space $H\times K^{\prime}$ under the group operation $G\times G\to G$ into the Hausdorff space $G$. Hence, $X^{H}$ is $K$-connected, since
$(X^{H})^{K^{\prime}}=X^{H}\cap X^{K^{\prime}}=X^{\langle H,K^{\prime}\rangle}=X^{HK^{\prime}},$
and $X$ is $G$-connected. Moreover, if $x_{0}\in X^{G}$, then $x_{0}\in(X^{H})^{K}$. Hence, by Lemma˜2.23, it is enough to show that $\mathrm{secat}_{K}(e_{X^{H}})\leq\mathrm{secat}_{G}(e_{X})$.
Corollary 2.25.
Let $G$ be a compact Hausdorff topological group, and let $X$ be a $G$-connected space with $X^{G}\neq\emptyset$. Then
@@
Now, as a consequence of Theorem˜2.16, we obtain an equivariant homotopy dimension-connectivity upper bound for equivariant LS category.
@@
Suppose $G$ is a compact Lie group. Suppose $X$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$ such that $X^{G}\neq\emptyset$. If there exists $s\geq 0$ such that $X^{H}$ is $s$-connected for all subgroups $H$ of $G$, then
@@
If $x_{0}\in X^{G}$, then $e_{X}\colon P_{x_{0}}X\to X$ is a $G$-fibration with fibre $\Omega X$, which also admits a $G$-action. Note that $(\Omega X)^{H}=\Omega(X^{H})$. Since $X^{H}$ is $s$-connected, the loop space $\Omega(X^{H})$ is (s1)(s-1)-connected. Hence, by Theorem˜2.16, we get
@@
As $X^{H}$ is $s$-connected, it follows that $X^{H}$ is path-connected. Hence, $X$ is $G$-connected, and the theorem follows by Lemma˜2.23. ∎
@@ 2.4. Equivariant and invariant topological complexity
We recall the concept of equivariant topological complexity introduced by Colman and Grant in [colmangranteqtc]. Let $X$ be a $G$-space. Observe that the free path space $PX$ admits a $G$-action via $(g\cdot\alpha)(t):=g\cdot\alpha(t)$. Similarly, the product space $X^{k}$ is a $G$-space with the diagonal action. The fibration
@@
In [A-D-S, Proposition 3.40], Sarkar and the authors of this paper provided a dimension-connectivity upper bound on the sequential equivariant topological complexity. We improve their result by establishing an equivariant homotopy dimension-connectivity upper bound. We omit the proof, as it is similar to the original and follows from the homotopy dimension-connectivity upper bound on the equivariant sectional category in Theorem˜2.16.
@@
Suppose $G$ is a compact Lie group. If $X$ is a $G$-$\mathrm{CW}$-complex of dimension at least 1 such that $X^{H}$ is $s$-connected for all subgroups $H\leq G$, then
$\mathrm{TC}_{k,G}(X)<\frac{k\ \mathrm{hdim}_{G}(X)+1}{s+1}+1.$
Example 2.29.
Consider $G=\mathbb{Z}_{2}$ acting on $X=S^{n}$ with $n\geq 1$, by reflection, given by multiplication by 1-1 in the last coordinate. Then $X^{G}=S^{n-1}$ and $X^{\{e\}}=S^{n}$ are (n2)(n-2)-connected and (n1)(n-1)-connected, respectively. Then, by Theorem˜2.26, it follows that
$\mathrm{cat}_{G}(S^{n})<\frac{n+1}{n-1}+1,\quad\text{for }n\geq 2,$
which implies $\mathrm{cat}_{G}(S^{n})\leq 2$ for $n\geq 3$. Hence, $\mathrm{cat}_{G}(S^{n})=2$ for $n\geq 3$, since $S^{n}$ is clearly not $G$-contractible. Moreover, by Theorem˜2.28, it follows that
$\mathrm{TC}_{k,G}(S^{n})<\frac{k\cdot n+1}{n-1}+1,\quad\text{for }n\geq 2$
which implies $\mathrm{TC}_{k,G}(S^{n})\leq k+1$ for $k\leq n-2$. Hence, it follows that $\mathrm{TC}_{k,G}(S^{n})=k+1$ for $k\leq n-2$, since
$\mathrm{TC}_{k}(S^{n-1})\leq\mathrm{TC}_{k,G}(S^{n})\quad\text{and}\quad\mathrm{TC}_{k}(S^{n})\leq\mathrm{TC}_{k,G}(S^{n})$
by [BaySarkarheqtc, Proposition 3.14 (2)], and
$\mathrm{TC}_{k}(S^{n})=\begin{cases}k,&\text{if }k\text{ is odd},\\ k+1,&\text{if }k\text{ is even},\end{cases}$
by [RUD2010, Section 4].
We note that this computation can be carried out more directly and efficiently as follows. Observe that $\mathrm{cat}_{G}(S^{n})=2$ for all $n\geq 2$, see [colmangranteqtc, Example 5.9]. Hence, it follows from [BaySarkarheqtc, Proposition 3.17] that $\mathrm{TC}_{k,G}(S^{n})\leq k+1$ for all $n\geq 2$ and all $k\geq 2$. Moreover, $\mathrm{TC}_{k,G}(S^{1})=\infty$, since $\mathrm{TC}_{k}(S^{0})\leq\mathrm{TC}_{k,G}(S^{1})$. Hence,
$\mathrm{TC}_{k,G}(S^{n})=\begin{cases}\infty,&\text{if }n=1,\\ k+1,&\text{if }n\geq 2.\end{cases}$
It is important to note that the equivariant topological complexity of $G$-spaces does not necessarily relate to the topological complexity of their orbit spaces. However, Lubawski and Marzantowicz [invarianttc] introduced a new notion of topological complexity for $G$-spaces, designed to facilitate such a comparison. We now present their definition and recall the corresponding result.
Suppose $X$ is a $G$-space. Let $\pi_{X}\colon X\to X/G$ denote the orbit map. Define
@@
Then the following diagram
[…]
@@
$\mathfrak{p}\colon PX\times_{X/G}PX\to X\times X,\quad(\gamma,\delta)\mapsto(\gamma(0),\delta(1)).$ (7)
It was shown in [invarianttc, Proposition 3.7] that the map $\mathfrak{p}$ is a $(G\times G)$-fibration.
@@
Moreover, if $X$ has only one orbit type, then
@@ 2.5. Clapp-Puppe invariant of Lusternik-Schnirelmann type
In this section, we recall the equivariant version of Clapp-Puppe invariant [clapp-puppe-ls-category], introduced by Lubawski and Marzantowicz in [invarianttc].
@@
Colman and Grant in [colmangranteqtc, Lemma 5.14] showed that for a $G$-invariant open subset $U$ of $X\times X$ the following are equivalent:
@@
Later, Lubawski and Marzantowicz in [invarianttc, Lemma 3.8] showed a similar result for invariant topological complexity. More precisely, for a $(G\times G)$-invariant open subset of $U$ of $X\times X$ the following are equivalent:
@@
In Section˜3 and Section˜4, we give analogous results for equivariant parametrized topological complexity and invariant parametrized topological complexity, respectively. We use these results to prove Theorem˜4.26.